OpenStudy (geekfromthefutur):

An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (geekfromthefutur):

@mathmale

OpenStudy (geekfromthefutur):

CAN ANYONE HELP ME!!!

OpenStudy (geekfromthefutur):

@saulglz211

OpenStudy (geekfromthefutur):

@igreen plz help me

OpenStudy (igreen):

Part A: You could say: I can graph the equation to find all the solutions to it. (Btw, remember me? I'm greenlegodude57, I just created a new account for a better username)

OpenStudy (geekfromthefutur):

oh ok, and holy crap whats up dude :D:D

OpenStudy (geekfromthefutur):

ohok so all i would need to say is i can graph the question to find my solution ??

OpenStudy (igreen):

Lol

OpenStudy (igreen):

To find ALL of the solutions. That's for Part A.

OpenStudy (geekfromthefutur):

oh ok thx bro can u help me with b and c lol

OpenStudy (igreen):

Yep

OpenStudy (igreen):

For Part B: Just say I looked on my graph and got these solutions: (0, 2),(1,11),(2, 20)

OpenStudy (geekfromthefutur):

jw how did u get those corresponds

OpenStudy (geekfromthefutur):

sorry i meant integers

OpenStudy (mathmale):

Part A: I'm assuming you're done with this. Your 9x - y = -2 becomes y=9x+2. Part B: all you have to do here is to choose three arbitrary x-values and to substitute them into y=9x+2, one at a time. You'll get three y-values. For example, if you choose x=2, then y=9(2)+2=18+2=20, and the corresponding point is (2,20). Find two more points in the same way (two more points that lie on this line). Part C: Before I discuss this, I'd like for you to look up the word "inconsistent" as it applies to algebraic equations. Perhqps look up "inconsistent systems of linear equations." Obviously we have to understand this terminology before we can come up with a valid answer to the question posed in Part C.